OpenStudy (anonymous):

x^4-y^4 ------- x^2+y^2

Parth (parthkohli):

What you will do here is that, you will factor the numerator and cancel something from the denominator.

Parth (parthkohli):

Keep in mind that \(x^4 - y^4 = (x^2)^2 - (y^2)^2\) which is somewhat a difference of two perfect squares.

OpenStudy (anonymous):

could you give me the sol and answer

Parth (parthkohli):

Not really... I am busy right now.

OpenStudy (anonymous):

ok thanks anyway

OpenStudy (anonymous):

Suppose X = x², Y = y² --> X² - Y² = ... @XDiesel Can you fill in where the .... is

OpenStudy (anonymous):

the what?

OpenStudy (anonymous):

i cant understand you

OpenStudy (anonymous):

X² - Y² because the variables are not the same. i guess

OpenStudy (mathmate):

You can use the standard identity: \[ a^2-b^2\ =\ (a+b)(a-b) \] and recognize that \[ x^4 - y^4 \] can be written as \[ (x^2)^2 - (y^2)^2 \]

OpenStudy (anonymous):

nice what is the final answer

OpenStudy (mathmate):

If you want to understand how to work out problems, this is the place. If you only want the answer, try: http://www.wolframalpha.com/ and type in your problem.

OpenStudy (anonymous):

thanks

OpenStudy (mathmate):

You're welcome! :)

OpenStudy (anonymous):

\[ \frac{x^4-y^4}{x^2+y^2}=\frac{\left(x^2-y^2\right) \left(x^2+y^2\right)}{x^2+y^2 }=x^2-y^2 \]

Parth (parthkohli):

Which could further be simplified, sir, to get \((x + y)(x - y)\)